### v1

# Causal Learning for Trustworthy Recommender Systems: A Survey

Abstract
Recommender Systems (RS) have significantly advanced online content discovery and personalized decision-making. However, emerging vulnerabilities in RS have catalyzed a paradigm shift towards Trustworthy RS (TRS). Despite numerous progress on TRS, most of them focus on data correlations while overlooking the fundamental causal nature in recommendation. This drawback hinders TRS from identifying the cause in addressing trustworthiness issues, leading to limited fairness, robustness, and explainability. To bridge this gap, causal learning emerges as a class of promising methods to augment TRS. These methods, grounded in reliable causality, excel in mitigating various biases and noises while offering insightful explanations for TRS. However, there lacks a timely survey in this vibrant area. This paper creates an overview of TRS from the perspective of causal learning. We begin by presenting the advantages and common procedures of Causality-oriented TRS (CTRS). Then, we identify potential trustworthiness challenges at each stage and link them to viable causal solutions, followed by a classification of CTRS methods. Finally, we discuss several future directions for advancing this field.
1 Introduction
Recommender Systems (RS), as one of the most successful applications of artificial intelligence, have revolutionized the way of our decision-making in the digital world. However, with their growing influence, concerns regarding the trustworthiness of RS have become increasingly prevalent. Inspecting the lifecycle of RS Wang et al. (2023), we identify several significant trustworthiness challenges across different stages, especially regarding fairness, robustness and explainability. They present pivotal yet challenging objectives for modern RS, extending beyond accuracy. Specifically, fairness requires unbiased decision-making for various users, items and providers. Robustness emphasizes the reliability of RS when confronted with noises or attacks. Explainability poses an urgent need for strong interpretability, advancing the understanding of RS mechanisms and results. Recently, due to its exceptional ability in identifying and utilizing causality, going beyond correlations, causal learning has been considered an effective and elegant tool to analyze and address Trustworthy RS (TRS) problems Gao et al. (2024); Li et al. (2023b). In this paper, we overview the trustworthiness challenges throughout the RS lifecycle, alongside a comprehensive exploration of causality-oriented solutions.
Figure 1: Illustrative comparison between conventional RS and CTRS. CTRS enhances robustness by learning underlying causality, whereas conventional RS may be misled by unreliable correlations.
1.0.1 Motivation: why Causality-oriented TRS (CTRS)?
Causality is more reliable than correlations. Conventional RS heavily rely on correlations learned from observed data. However, correlations are not always causally correct and spurious correlations can be notably misleading, as exemplified in Figure 1. In contrast, CTRS methods He et al. (2023) are adept at identifying and modeling genuine causal relations for enhanced robustness. Moreover, another persistent issue in RS is the non-random nature of missing data Wang et al. (2019), as user-item interactions are typically preference-driven, rather than random. By leveraging advanced causal techniques (e.g., propensity reweighting Li et al. (2023a)), the impacts of this issue can be effectively mitigated, paving the way for more accurate and reliable recommendations.
Causal learning offers theoretical instructions for identifying and eliminating biases. Causal learning provides various theoretically grounded approaches to analyze the bias in RS via straightforward yet insightful causal models. Upon targeting the true cause of biases, a wide range of causal techniques (e.g., structure adjustment and counterfactual inference) can be applied to measure and adjust the underlying cause-and-effect relationships. Benefiting from this, a significant enhancement in the fairness of RS can be achieved.
Causal learning helps provide straightforward and effective explanation. Conventional RS models, which are commonly considered black-box models, can hardly offer clues about the reason for recommendations. CTRS, on the other hand, empowered by path discovery and counterfactual inference, possess superior interpretability. These methods Xian et al. (2019); Yu et al. (2023) not only shed light on the recommendation mechanism but also indicate the rationale behind the results, thereby providing more transparent and comprehensible frameworks for both users and developers.
1.0.2 How does causal learning help solve TRS problems?
Causal learning shows promising efficacy in understanding the cause-and-effect relationships of variables in a TRS problem. Typical CTRS methods include two major steps, i.e., causal formulation and causal inference. The former step aims to accurately delineate causal relationships within the system. This is usually achieved by the Structural Causal Model (SCM) Pearl (2009). Formally, it is defined as $\mathcal{G}=(\mathcal{V},\mathcal{E})$, where $\mathcal{V}$ is a set of nodes, representing variables associated with the problem, and $\mathcal{E}$ is a set of edges, illustrating underlying causal relationships. Following such established causal models, either through the process of causal discovery or leveraging prior knowledge, the root causes of problems related to trustworthiness can be clearly identified. Subsequently, advanced causal inference tools can be applied to intervene in causality and mitigate effects of these problems.
Despite several surveys on TRS Wang et al. (2023); Ge et al. (2022); Fan et al. (2022), as well as causal learning related RS Gao et al. (2022); Luo et al. (2023); Zhu et al. (2023); Xu et al. (2023b); Saito and Joachims (2021); Wu et al. (2022b), they lack a specific focus on CTRS. Surveys on TRS typically concentrate on data-driven methods while these methods are prone to issues of spurious correlations, biases and limited interpretability, as discussed earlier. Meanwhile, surveys on causal learning for RS often focus on technical causal inference tools applied in RS, neglecting the connection with trustworthiness challenges. Moreover, most of them overlook the importance of ensuring trustworthiness across different stages. To fill these gaps, we provide a structured survey to examine trustworthiness challenges throughout the RS lifecycle and align them with causal solutions.
Contributions. The main contributions of this survey are summarized as follows:
  • It outlines trustworthiness challenges in the lifecycle of RS and aligns them with viable causal solutions, highlighting the significance and necessity of causal learning in addressing TRS problems.
  • It summarizes the current progress of CTRS, aiming to hold practical values in guiding the development of unified trustworthy RS platforms.
  • It offers valuable insights into future directions and perspectives for the future advancement of CTRS.
RS stage
Aspect
Sub-aspect
Trustworthiness challenge
Causal solution
Example
Data preparation
Fairness
User side
User attribute bias
Counterfactual inference
[1]
Population bias
Counterfactual augmentation
Item side
Item attribute bias
Counterfactual inference
[1]
Popularity bias
Structure adjustment
[2], [3]
Counterfactual inference
[4]
Position bias
Propensity reweighting
[5]
Robustness
Non-randomness
Selection bias
Propensity reweighting
[6], [7], [8]
Corruption
Poisoning attacks
Counterfactual augmentation
Representation learning
Fairness
Implicit intention
Conformity bias
Causal disentanglement
[9]
Robustness
Noise
Spurious correlations
Structure adjustment
[10]
Counterfactual augmentation
[11], [12]
Noisy feedback
Counterfactual inference
[13]
Recommendation generation
Fairness
Model side
Bias amplification
Structure adjustment
[14]
Filter bubble bias
Counterfactual inference
[15], [16]
Explainability
Model agnostic
Unexplainable results
Counterfactual inference
[17], [18], [19]
Model specific
Black-box mechanism
Path discovery
[20]
Evaluation
Robustness
Model effect
Uplift evaluation
Uplift effect estimation
[21]
[1]Li et al. (2021); [2]Zhang et al. (2021c); [3]Gupta et al. (2021); [4]Wei et al. (2021); [5]Joachims et al. (2018); [6]Zhang et al. (2020b); [7]Li et al. (2023b); [8]Li et al. (2023a); [9]Zheng et al. (2021); [10]He et al. (2023); [11]Mu et al. (2022); [12]Xu et al. (2023a); [13]Zhang et al. (2021b); [14]Wang et al. (2021b); [15]Wang et al. (2022b); [16]Gao et al. (2024); [17]Tan et al. (2021); [18]Chen et al. (2022); [19]Yu et al. (2023); [20]Xian et al. (2019); [21]Sato et al. (2019)
Table 1: Trustworthiness challenges in the lifecycle of RS and corresponding causal learning solutions.
2 Trustworthiness Challenges in RS Lifecycle
A typical recommender system is mainly composed of four successive and inter-connected stages, i.e., data preparation, representation learning, recommendation generation and evaluation. Each stage plays a fundamental role in the lifecycle of RS Wang et al. (2023) and is confronted with its own trustworthiness challenges. We highlight that the overall trustworthiness of RS essentially hinges on the reliability of each stage. Only when each stage is deemed trustworthy can we develop a unified trustworthy RS. In this section, we delve into the trustworthiness challenges inherent at each stage, examining them through the lens of causal learning. Table 1 summarizes these challenges across the various stages and illustrates their alignment with causal solutions, offering a structured overview of the landscape.
2.1 Challenges in Data Preparation
As the first stage of RS, data preparation involves gathering, organizing and transforming input data (e.g., user-item interactions and side information). However, raw data collected from complex cyberspace can be insufficiently trustworthy and commonly carry various biases (e.g., popularity bias), noises (e.g., fake ratings/reviews from online water army) and even malicious attacks. This results in substantial challenges to the overall trustworthiness of the system, especially in ensuring fairness and robustness in data preparation.
2.1.1 Fairness Challenges in Data Preparation
During data preparation, a primary fairness challenge lies in how to mitigate the pre-existing biases Li et al. (2023c) in raw data. Neglecting such inherent biases will pose serious fairness risks to those RS models trained on these biased data, potentially harming the interests of stakeholders. According to different subjects of fairness concerns, there exist two major types of pre-existing biases, i.e., user-side Li et al. (2021) and item-side bias Zhang et al. (2021c).
User-side bias refers to unfair recommendations for different groups of users. The first prevalent bias in this category, i.e., user attribute bias, refers to unequal treatment of users based on their sensitive attributes (e.g., gender, race, and educational background). For example, a biased job recommender may make discriminatory decisions based on the gender of users Mansoury et al. (2020a). Another critical user-side bias, namely population bias, stems from the unbalanced distribution of users. Typically, heavy users contribute a significantly larger volume of interaction data than casual users, leading to a highly unbalanced dataset. Consequently, RS models trained on such data tend to favor heavy users and perform noticeably better Wang et al. (2022a); Wu et al. (2022a) for them than for casual users.
Additionally, in a trustworthy recommendation ecosystem Wang et al. (2023), mitigating item-side bias helps to safeguard the interests of items and their providers. Similar to users, items may suffer from attribute bias because of certain properties. However, beyond general attributes, such as brands and production places, the popularity of an item can markedly affect the recommendation results, giving rise to popularity bias Zhang et al. (2021c). Popular items, benefiting from their fame, naturally draw more attention than long-tail products. If RS models capture and rely on these patterns, they may blindly prefer to recommend popular items without adequately considering the user-item alignment. Another classic item-side bias, namely position bias Joachims et al. (2018), is based on the common observation Joachims et al. (2007) that users tend to pay more attention and are more likely to interact with items at the top of a recommendation list, regardless of their actual relevance or quality. Therefore, such biased interaction data may mislead newly trained models, resulting in unfairness to other items.
Several efforts Wu et al. (2021) have been made in recent years to address the aforementioned pre-existing biases. For instance, Abdollahpouri et al. Abdollahpouri et al. (2017) developed a regularization-based framework to achieve a balance between accuracy and fairness objectives. However, these data-driven methods, relying on patterns observed in data, often fall short in understanding the mechanism of how bias affects interactions. This significantly limits their efficacy in debiasing within practical applications. To bridge this gap, several causality-oriented methods have been proposed, focusing on analyzing and measuring the causal relationships within a biased RS. For user-side bias, counterfactual reasoning stands out as an elegant solution, which can not only achieve counterfactual fairness Li et al. (2021) for individuals by reducing the impacts of sensitive attributes, but also can help to generate reasonable samples Mu et al. (2022) for minority users. For popularity bias, existing work commonly formulates the problem within the framework of SCM and performs structure adjustment Zhang et al. (2021c) or counterfactual inference Wei et al. (2021) for debiasing. For position bias, Joachims et al. Joachims et al. (2018) devised a propensity-weighted model to reduce the bias caused by positions in ranking. The details of causal learning methods will be introduced in Section 3.
2.1.2 Robustness Challenges in Data Preparation
Since raw data are usually collected from uncontrollable environments, it is pivotal to handle the robustness challenges stemming from non-randomness Li et al. (2023a) and corruption Zhang et al. (2022) during data preparation.
Due to the inherent high sparsity of user-item interactions, most existing work assumes that interactions are Missing At Random (MAR) for modeling simplicity and statistical convenience. However, this assumption does not always hold in practice. In cases where samples are Missing Not At Random (MNAR), selection bias Li et al. (2023b) becomes a significant issue. It refers to the distortion that data used to train RS models are not representative of the entire range of user preferences. To overcome this challenge, earlier work Hernández-Lobato et al. (2014) managed to model the generative and missing process of data, but this led to high computational costs. Recently, propensity reweighting tools have been widely adopted to solve selection bias in RS. These methods offer higher efficiency Schnabel et al. (2016) and flexibility to integrate with various recommendation frameworks Zhang et al. (2020b); Li et al. (2023a).
Another critical robustness challenge in data preparation is how to defend against poisoning attacks Zhang et al. (2022). These attacks involve the insertion of fake users into training data, with the intent to manipulate the outcomes of RS models for malicious purposes. To counteract such threats, two representative defense strategies are primarily adopted, i.e., detection-based methods Zhang et al. (2020a) and adversarial training Tang et al. (2020). Detection-based methods focus on developing an accurate classifier to identify anomalies and spurious samples. Adversarial training, on the other hand, aims to enhance model robustness by introducing generated perturbations during training. However, these defense methods are essentially data-driven. There remains an imperative need to explore causality-oriented defense methods, pursuing a deeper understanding of how attacks manipulate RS results. Useful causal learning tools include counterfactual augmentation for adversarial training and causal effect inference for causal measurement.
2.2 Challenges in Representation Learning
Representation learning plays a fundamental role in bridging the stages of data preparation and recommendation generation. Although numerous advanced representation learning methods have been developed in RS, such as Recurrent Neural Networks (RNN), Graph Neural Networks (GNN) Ma et al. (2020) and transformer-based models Sun et al. (2019), they commonly overlook potential risks regarding fairness and robustness.
2.2.1 Fairness Challenges in Representation Learning
A major fairness challenge during representation learning is how to address conformity bias Zheng et al. (2021). This bias is closely related to popularity bias, but it specifically focuses on the perspective of user intentions. Generally speaking, a user-item interaction is usually caused by two factors: users’ genuine interests in the item and their conformity to the item’s popularity. An unbiased model should focus on learning users’ true interests rather than being misled by their conformity. However, it is intractable to clearly separate conformity intention from user representation, especially relying solely on data-driven approaches. To tackle this challenge, Zheng et al. Zheng et al. (2021) introduced a causality-driven framework, dubbed DICE, to learn representations with disentangled intentions. It presents a simple SCM model to depict the causal relationships between user implicit intentions and the resulting interactions. Leveraging the principle of the colliding effect Pearl (2009), DICE is adept at learning disentangled representations from cause-specific data.
2.2.2 Robustness Challenges in Representation Learning
Apart from explicit robustness risks in data, underlying noises can also significantly hinder the effectiveness of representation learning. Here, we mainly focus on two types of noises, i.e., spurious correlations He et al. (2023) and noisy feedback Zhang et al. (2021b).
Spurious correlation refers to a correlation that is statistically strong but in fact is causally incorrect. To clarify this concept, consider a straightforward example. It is a well-established fact that an item’s quality influences both its price and the users’ preference for it. Commonly, one can observe that higher-quality items are more expensive and more appealing to users. However, if the factor of quality is neglected, an RS may falsely imply that user preference increases along with the price. Such spurious correlations can be largely misleading. Unfortunately, data-driven methods, which typically rely on correlations or associations among samples, can hardly identify and eliminate such noises. In contrast, causality-oriented methods excel in tackling this issue by modeling and adjusting underlying causal relationships. For instance, recognizing that spurious correlations often stem from confounders, a study by Zhu et al. (2022) performs structure adjustment to mitigate corresponding confounder effects. Another representative research by Mu et al. (2022) introduces a counterfactual-based method to bolster model robustness against spurious correlations by generating high-quality counterfactual interactions.
Noisy feedback manifests in two ways: either as positive interactions made by users who in the end turned out to be unsatisfied Wang et al. (2021a, c), or as delayed interactions that are actually satisfied Zhang et al. (2021b). Such noisy feedback eventually limits the ability of an RS to capture users’ genuine preferences. To defend against this type of noise, Zhang et al. Zhang et al. (2021b) devised a counterfactual modifier in a reinforcement learning framework for streaming recommendation. It adopts importance sampling Yasui et al. (2020) and modifies the weight of reward in a counterfactual world to alleviate the impact of delayed feedback.
2.3 Challenges in Recommendation Generation
Upon obtaining informative representations from input data, RS models are devised to generate accurate suggestions that are tailored to individual preferences. Despite the promising performance that modern RS models Lin et al. (2022) achieved, there still exist several challenges regarding fairness and explainability due to their intrinsic mechanisms.
2.3.1 Fairness Challenges in Recommendation Generation
As RS models are commonly trained dynamically through successive feedback loops Mansoury et al. (2020b), biases existing in results and user feedback will be continuously amplified, known as bias amplification. By analyzing underlying causal relationships, Wang et al. Wang et al. (2021b) attribute bias amplification to the confounding effect arising from users’ imbalanced historical distribution. Hence, they perform backdoor adjustment Pearl (2009) to remove causal paths from confounders, thus mitigating the impact of this issue.
Additionally, feedback loops in RS often reinforce users’ existing preferences, creating an echo chamber that isolates them from diverse contents, known as the filter bubble bias Gao et al. (2024). To counter this challenge, a framework based on counterfactual inference Wang et al. (2022b) has been developed. This approach allows users to determine whether actions are needed to modify user embeddings through counterfactual modification. By doing so, it facilitates access to a wider and more diverse range of contents, effectively helping users to break out of filter bubbles.
2.3.2 Explainability Challenges in Recommendation Generation
The advancement of machine learning-based, especially deep learning-based techniques, has greatly enhanced the capability of RS models over the past decades. However, the inherent drawback of lacking interpretability has garnered significant concerns in recent years. With the introduction of complex black-box modules, it is challenging to explain how and why the recommendations are generated. This limitation not only impedes researchers in comprehensively understanding and improving models, but also undermines stakeholders’ trust in RS, especially in some safety-critical applications. Therefore, effective explainers are required for providing insightful explanations for both recommendation mechanisms and results.
Despite several attempts Zhang and Chen (2020) to generate explanations based on tags Vig et al. (2009) or visual attentions Chen et al. (2019), they are not convincing enough and often struggle to explain the mechanism of RS models. To bridge this gap, causality has been explored in RS to generate insightful explanations. For instance, a reinforcement learning model Xian et al. (2019) is designed to discover reasonable causal paths as explanations for how recommendations are generated. However, such path-based explanations also raise concerns regarding privacy leakage. Another representative work Tan et al. (2021) introduces a model-agnostic explainer based on counterfactual inference. This explainer aims to find straightforward explanations by searching for a set of changes on feature vectors, which is the least complex but strong enough to alter recommendation outcomes.
2.4 Challenges in Evaluation
Current accuracy-oriented evaluation metrics usually concentrate on the interaction rate of the recommendation results. However, considering the fact that users are free to interact with items they want regardless of the results, an RS should focus more on the uplift of interactions that are actually brought by recommendations. This poses a significant challenge, as it is intractable to observe the interaction outcomes with and without recommendations at the same time. To address this issue, Sato et al. Sato et al. (2019) proposed a causal inference framework for RS. It leverages additional interaction data of an existing model to estimate the uplift effects of a new model. Moreover, it presents a novel optimization algorithm to promote the uplift of interactions during training.
Moving beyond accuracy, TRS encompass a variety of trustworthiness objectives, posing the necessity of specialized evaluation metrics. In particular, fairness metrics Li et al. (2023c) typically measure group or individual disparities. For instance, Wang et al. Wang et al. (2022b) employed an isolation index to measure the segregation across different user groups and to assess the severity of filter bubbles. In evaluating robustness, most works Li et al. (2023a) test RS accuracy under varying levels of robustness challenges. In addition, Li et al. Li et al. (2023b) proposed a novel metric to assess the balance of learned propensity weights when dealing with selection bias. In evaluating explainability, user reviews Yu et al. (2023) are usually collected to assess the consistency with generated explanations. Moreover, Tan et al. Tan et al. (2021) introduced two novel metrics, based on the concepts of necessity and sufficiency, to provide a thorough evaluation of explanation quality.
However, it should be pointed out that only a few attempts Sato et al. (2019) have been made to devise evaluations from a causal learning perspective. It is an important yet underexplored area in CTRS and it deserves wider research attention in the future. Moreover, we keep in mind that the trustworthiness of RS is a human-centered consideration. Therefore, apart from the technical side, evaluations on user perception and ethical side Wang et al. (2023) are equally vital. This therefore may require user studies to assess the subjective perceptions of fairness, robustness, explainability and even potential societal impacts.
Figure 2: Classification of CTRS methods.
Approach Applicable challenge Pros Cons Example
Causal discovery
Causal disentanglement
Conformity bias
An effective way to separate interaction data based on their primary causes
Relies on prior knowledge of causal models and is hard to scale to complex scenarios
[1]
Path discovery
Black-box mechanism
Straightforward to illustrate the process of recommendations
Potential risks of privacy leakage
[2]
Causal effect inference
Propensity reweighting
Position bias, selection bias
Proven to be statistically unbiased
Hard to accurately and robustly estimate the propensity weights
[3], [4], [5], [6]
Structure adjustment
Popularity bias, spurious correlations, bias amplification
Effective in deconfounding and estimating accurate causal effects
Relies on given causal models
[7], [8], [9], [10]
Uplift effect estimation
Uplift evaluation
Insightful to show the actual promotion brought by recommendations
Additional training data are required, which can be difficult to obtain in real-world applications
[11]
Counterfactual reasoning
Counterfactual augmentation
Population bias, poisoning attacks, spurious correlations
Excels in generating sufficient samples for training
Lack of assurance of the quality of generated samples
[12], [13]
Counterfactual inference
Attribute bias, popularity bias, noisy feedback, filter bubble bias, unexplainable results
A feasible way to explore causal effects of hypothetical interventions
Assumptions involved in computation may not always hold in practice
[14], [15], [16], [17], [18], [19], [20], [21]
[1]Zheng et al. (2021); [2]Xian et al. (2019); [3]Joachims et al. (2018); [4]Zhang et al. (2020b); [5]Li et al. (2023b); [6]Li et al. (2023a); [7]Zhang et al. (2021c); [8]Gupta et al. (2021); [9]He et al. (2023); [10]Wang et al. (2021b); [11]Sato et al. (2019); [12]Mu et al. (2022); [13]Xu et al. (2023a); [14]Li et al. (2021); [15]Wei et al. (2021); [16]Zhang et al. (2021b); [17]Wang et al. (2022b); [18]Gao et al. (2024); [19]Tan et al. (2021); [20]Chen et al. (2022); [21]Yu et al. (2023);
Table 2: A comparison of different classes of CTRS methods.
3 Causality-oriented TRS Methods
Recently, causal learning has stood out as an appealing solution to addressing trustworthiness challenges in the construction of TRS. We first present a taxonomy of existing CTRS methods, as illustrated in Figure 2, organized by different causal learning techniques upon which they are built. Then, we discuss their technical details and compare different classes of CTRS methods in Table 2, regarding their applicable challenges, pros and cons, and representative works.
3.1 Causal Discovery based Methods
Causal discovery is a crucial aspect of causal learning, focusing on identifying causal relationships from data. Unlike traditional data-driven methods that emphasize correlation, causal discovery based methods seek to uncover the causality among variables and formulate it in causal models. This is particularly valuable in RS, where understanding the causal relationships between each variable (e.g., users’ intention or items’ popularity) and the outcome (e.g., user-item interactions) is essential yet intractable. For different types of RS data, two representative causal discovery based methods have been researched: Causal Disentanglement based RS (CDRS) and Path Discovery based RS (PDRS).
3.1.1 Causal Disentanglement based RS (CDRS)
Interaction data are commonly influenced by various factors and underlying causes. For addressing fairness issues, it is essential to learn representations that remain unaffected by bias factors, such as user conformity. CDRS methods Zheng et al. (2021) address this by leveraging the colliding effect to distill cause-specific data from the original training dataset. This enables the learning of causally disentangled representations from the refined data, effectively reducing the impact of conformity bias, a common challenge in traditional RS. However, CDRS methods heavily rely on prior knowledge of causal relationships. Meanwhile, the complexity of disentanglement increases as more factors are involved.
3.1.2 Path Discovery based RS (PDRS)
Recent years have witnessed an upsurge in leveraging knowledge graphs Yang et al. (2022) for accurate recommendations. To explore the causality during knowledge graph reasoning and to enhance the understanding of the recommendation mechanism, PDRS Xian et al. (2019) trains a reinforcement learning agent to learn the navigation to suitable items for each user. An efficient path reasoning algorithm is devised to sample possible paths, which can be used as straightforward explanations of how recommendations are generated.
3.2 Causal effect inference based Methods
Based on the discovered or given causality in an RS, causal effect inference aims to understand and quantify the effect of one variable on another. Generally, causal effect inference based methods can be grouped into Propensity Reweighting based RS (PRRS), Structure Adjustment based RS (SARS) and Uplift Effect Estimation based RS (UEERS).
3.2.1 Propensity Reweighting based RS (PRRS)
Propensity reweighting is a set of fundamental techniques in causal effect inference, designed to adjust the biased observations in scenarios where a random experimental assignment is impractical. This makes it particularly useful in addressing position bias Joachims et al. (2018) and selection bias Li et al. (2023b) in RS. PRRS methods usually involve three key steps, i.e., propensity weight estimation, weight assignment and unbiased recommendation. In detail, propensity weight estimation aims to calculate the probability of an item receiving treatment (e.g., interaction). For example, in the problem of post-click conversion rate estimation Zhang et al. (2020b), the propensity weight is the click-through rate, which can be predicted through a logistic regression model. The inverse of the propensity weight can be then applied to the loss target, following the idea of Inverse Propensity Scoring (IPS) Hirano et al. (2003). Finally, these adjustments help mitigate biases, leading to more robust and fair recommendations. However, PRRS methods commonly face challenges with accuracy in weight estimation and high variance due to inverse weighting. Fortunately, recent advancements, such as IPS-V2 Li et al. (2023b) and doubly robust model Wang et al. (2019), have been developed to tackle these issues.
3.2.2 Structure Adjustment based RS (SARS)
In a given system where the underlying causal relationships are modeled by an SCM, structure adjustment emerges as a pivotal technique to accurately estimate causal effects, particularly by mitigating the influence of confounders. Recognizing that confounders are often the primary sources of various biases and noises, structure adjustment has gained great attention in building TRS. Depending on the variety of confounders, SARS methods usually adopt two types of adjustment tools. The first one is known as backdoor adjustment Zhang et al. (2021c), which leverages the backdoor criterion Pearl (2009) to identify and adjust observed confounders. Frontdoor adjustment Xu et al. (2023c), as another common technique, shows particular value in cases with unobserved confounders. Following the frontdoor criterion Pearl (2009), it finds a mediator to help in estimating the causal effects when direct adjustments for confounders are not possible.
3.2.3 Uplift Effect Estimation based RS (UEERS)
Uplift in RS refers to the increase in user actions that are genuinely caused by recommendation outcomes. Normally, assessing these uplift effects requires knowing whether the user would interact with the item if it were recommended or not, which is, however, unfeasible to observe in practice. To handle this, Sato et al. Sato et al. (2019) present an uplift evaluation based on additional interaction data observed from an existing RS model. It estimates the uplift effect of a newly deployed model by calculating the average treatment effect and leverages propensity reweighting tools to mitigate the bias arising from the non-randomness of observed data. Despite its effectiveness, the requirement for additional data in UEERS can be costly and challenging, leading to limited practicality.
3.3 Counterfactual Reasoning based Methods
Moving beyond causal effect inference, counterfactual reasoning focuses on discovering effects of interventions in hypothetical scenarios to address the “what if” questions. Derived from this concept, two influential tools have been widely used in CTRS: Counterfactual Augmentation based RS (CARS) and Counterfactual Inference based RS (CIRS).
3.3.1 Counterfactual Augmentation based RS (CARS)
Counterfactual augmentation refers to the process of enriching data within counterfactual scenarios. This usually involves intervened synthetic data points that represent plausible alternative outcomes under different conditions. For example, Zhang et al. Zhang et al. (2021a) introduced a module for this purpose to address missing data issues. It generates positive and negative samples by replacing non-essential features and essential features, respectively. By incorporating counterfactual samples, these CARS methods can be better equipped to handle diverse and unseen biases, noises and attacks, leading to reliable recommendations in practice. However, there remains a need for effective strategies to assess and ensure the quality of counterfactual samples.
3.3.2 Counterfactual Inference based RS (CIRS)
As another influential tool, counterfactual inference is commonly used to evaluate and leverage causal relationships, aiming to answer the “what if” question. These questions vary across different problems. For instance, to handle the popularity bias in RS, Wei et al. Wei et al. (2021) investigated the question: “what the ranking score would be if the model only uses item properties”. By measuring the causal effects of two alternative situations, they identified that the key to mitigating popularity bias lies in reducing the direct effect of a counterfactual RS, which assigns ranking scores without considering the user-item match. Accordingly, an effective and compatible CIRS framework is devised for a wide range of existing models to alleviate the impacts of popularity bias.
4 Open Problems and Future Directions
In this section, we highlight several open problems and directions for future research.
Unresolved trustworthiness challenges. As shown in Table 1, there exist several trustworthiness challenges that are critical but remain unresolved by CTRS methods, such as population bias and poisoning attacks. Although we have presented some preliminary ideas for addressing these challenges through lens of causal learning, there is a clear need for more in-depth research to develop comprehensive solutions.
Beyond causal prior knowledge. Most existing CTRS methods He et al. (2023); Zheng et al. (2021) rely on handcrafted causal analysis and pre-defined causal models. This significantly limits their practical effectiveness, especially in scenarios with intricate causal relationships. In causal learning, causal discovery is a fundamental topic, where various advanced techniques have been researched in recent years, including SAT Hyttinen et al. (2014) and $\rm A^{*}$ based methods Lu et al. (2021). However, they have not received much attention in CTRS. Hence, there is a growing need to incorporate advanced causal discovery techniques into CTRS for enhanced practicality and applicability.
Causal-oriented trustworthy evaluations. There has been limited research on evaluating RS performance from causal learning perspectives Sato et al. (2019). Developing advanced causal-oriented evaluations helps to bridge the gap between causal theory and practical RS assessment. Moreover, existing trustworthiness evaluations primarily focus on assessing recommendation results, while ensuring trustworthiness across all stages of the RS lifecycle is crucial. Hence, multi-stage trustworthy evaluations are in great demand.
Towards more generalized situations. When delving into theoretical details of causal effect inference and counterfactual reasoning based methods, one can find that they essentially rely on several fundamental assumptions (e.g., consistency assumption) Pearl (2009). These assumptions, however, may not always hold in practical settings. Therefore, it is critical to develop and introduce advanced causal frameworks that offer enhanced generalization capabilities, thereby accommodating a wider range of real-world scenarios.
5 Conclusions
Recommender systems have become an integral part of our daily lives, playing a crucial role in shaping user experiences. Specifically, there has been an increasing focus on the trustworthiness of RS, where a variety of Causal-oriented Trustworthy RS (CTRS) methods emerge using the advancement of causal learning to address fairness, robustness and explainability challenges in RS. However, no systematic reviews are available. This survey presents an overview of CTRS methods. We not only identify major trustworthiness challenges at each stage of the RS lifecycle but also clarify the necessity of causal learning, providing valuable insights into the practical development and enhancement of CTRS. In addition, we have collected and organized the representative open-source CTRS algorithms and datasets on https://github.com/JinLi-i/CTRS_resource.